OpenStudy (anonymous):

Help please! Given the function f(x) = log3(x + 4), find the value of f^−1(3).

OpenStudy (anonymous):

f−1(3) = 7 f−1(3) = 12 f−1(3) = 23 f−1(3) = 31

OpenStudy (anonymous):

c

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw